Solve for x:
2x – 3 = -4

Answer by tutorcecilia(2144)

Solve for x:
2x – 3 = -4 [solve for x]
.
2x – 3 + 3 = -4 + 3
2x = -1
2x/2 = -1/2
x = -1/2
.
Check by pluggin-in (x=-1/2) back into the original equation.